# A method to derive material-specific spin-bath model descriptions of materials displaying prevalent spin physics (for simulation on NISQ devices)

### v2

@@ III.5 Molecular Chromium bromide
We observe that the ground state energies of the Hamiltonians $\mathcal{P}H\mathcal{P}$, $\tilde{H}$, $\mathcal{P}\tilde{H}\mathcal{P}$ are within a fraction of a Hartree from the true ground state energy $E_{0,H}$ of the original Hamiltonian $H$. The respective energy differences are $\Delta E_{0,\tilde{H}}=(E_{0,H}-E_{0,\tilde{H}})\simeq 1\times 10^{-4}\,\mathrm{Hartree}<(E_{0,H}-E_{0,\mathcal{P}\tilde{H}\mathcal{P}})\simeq 7\times 10^{-3}\,\mathrm{Hartree}<(E_{0,H}-E_{0,\mathcal{P}H\mathcal{P}})\simeq 8\times 10^{-3}\,\mathrm{Hartree}$. In contrast, we find a significantly larger energy discrepancy $(E_{0,H}-E_{0,\mathcal{P}_{0}H\mathcal{P}_{0}})\simeq 2\times 10^{-1}\,\mathrm{Hartree}$ for the effective Hamiltonian $\mathcal{P}_{0}H\mathcal{P}_{0}$.
@@
At first glance, we observe significant energy differences between the full Hamiltonian $H$ and the effective Hamiltonians $\mathcal{P}H\mathcal{P}$ and $\mathcal{P}\tilde{H}\mathcal{P}$ for the excited states, while the transformed Hamiltonian $\tilde{H}$ appears to remain a reasonable approximation to the original Hamiltonian $H$ with $(E_{n,H}-E_{n,\tilde{H}})\leq 2\times 10^{-3}\,\mathrm{Hartree}$. We find that the $n=9$ lowest excited states correspond to an average local parity $\overline{\langle P_{q\in\mathcal{S}}\rangle}_{n}=\frac{1}{3}\sum_{q\in\mathcal{S}}\langle P_{q}\rangle_{n}\geq-\frac{1}{3}$. This implies the presence of just one single spin degree of freedom instead of the three, such that the effective three-spin-system Hamiltonians $\mathcal{P}H\mathcal{P}$ and $\mathcal{P}\tilde{H}\mathcal{P}$ are not sensible descriptions of the physics of the excited states with significant charge fluctuations. If we reduce the spectrum to eigenstates of $H$ with three spin-like orbitals, i.e. average local parity $\overline{\langle P_{q\in\mathcal{S}}\rangle}_{n_{\mathcal{S}}}\simeq-1$, a different picture emerges. This reduced spectrum is shown in Fig. 8 (b). The effective Hamiltonians $\mathcal{P}H\mathcal{P}$ and $\mathcal{P}\tilde{H}\mathcal{P}$ appear to capture this reduced spectrum well and we observe a maximum energy discrepancy $\Delta E_{n_{\mathcal{S}}}=2\times 10^{-2}\,\mathrm{Hartree}$. The effective Hamiltonian $\mathcal{P}_{0}H\mathcal{P}_{0}$ remains a bad approximation throughout this reduced energy spectrum. This highlights the necessity of a suitable orbital basis for the derivation of effective Hamiltonians describing the spin physics. In the parity-optimized orbital basis, for the given example of $\mathrm{Cr}\mathrm{Br}^{3+}_{6}$, the terms coupling to the charge degree of freedom of the electron densities in the spin-like orbitals $\phi_{q\in\mathcal{S}}$ are suppressed to the extent that the simple effective Hamiltonian $\mathcal{P}H\mathcal{P}$ already becomes a good description of the spin excitation spectrum. The subsequent Schrieffer-Wolff transformation and utilization of the corresponding effective Hamiltonian $\mathcal{P}\tilde{H}\mathcal{P}$ yields insignificant improvements. This emphasizes the importance of using a suitable single particle basis. In constrast, we observe no such drastic suppression of the undesirable coupling terms for the basis of the CASSCF canonical molecular orbitals.
@@
We have shown that our optimization procedure is able to correctly identify the $N_{\phi_{q}}=3$ spin-like orbitals such that a subset of the energy spectrum of $\mathrm{Cr}\mathrm{Br}^{3+}_{6}$ can be accurately reproduced by a system of three spin degrees of freedom coupled to a fermionic environment. We further observe that our Schrieffer-Wolff transformation approach yields a transformed Hamiltonian $\tilde{H}$ that, if projected to the Hilbert space representing three spins in a fermionic environment, at least equals but typically betters the description given by the original Hamiltonian $H$ projected to the same subspace. We validate that the local parity of the spin-like basis orbitals can be regarded as a good indicator of the accuracy of the effective spin-bath Hamiltonian description $\mathcal{P}\tilde{H}\mathcal{P}$ derived with our method.
@@ IV Conclusion
Effective model Hamiltonians provide enormous value for the investigation of individual aspects of complex systems. Here, we have presented a procedure for the automatic derivation of model Hamiltonians of coupled spin degrees of freedom embedded in a fermionic environment. These Hamiltonians provide a means to enhance the study of the dominant low temperature spin physics of complex real materials. We have shown that to investigate the spin physics of such materials properly, one first needs to operate within a suitable single particle orbital basis, where a subset of the basis orbitals become realizations of spin degrees of freedom at low energies. We introduced the local parity as a sensible metric for the spin-like behavior of electron densities in the basis orbitals. We have proven that the natural orbital basis does not generally correspond to the basis that features the spin-like orbitals. We have provided an example toy model for which the natural orbital basis is significantly worse than the initial basis for the study of spin physics. We also studied a well-known singlet-diradical molecule, namely closed configuration para-benzyne, and demonstrated that the two well-known non-trivial spin-like orbitals are correctly determined by our local parity optimization procedure. We find that the optimization of the local parity yields an orbital basis that typically results in a many particle basis of the Hilbert space with small coupling and significant energy gap between the different subspaces. Our extended Schrieffer-Wolff transformation method allows to also transform generic Hamiltonians with complicated interaction terms beyond the density-density interactions of the toy models that the Schrieffer-Wolff transformation has so far mostly been applied to. There is also no need for the block-diagonal part $H_{0}$ to be easily diagonalizable, because no calculation of the spectrum of $H_{0}$ is required in Schrieffer-Wolff transformation our method. We have shown that our Schrieffer-Wolff transformation method reproduces established results for the well-known single impurity Anderson model and the disordered Fermi-Hubbard model Hamiltonians. For these models we corroborate that the local parity of the spin-like orbitals is a good predictor for the quality of the effective spin-bath Hamiltonian as the description of the low energy spin dynamics. Lastly, we have applied the full spin-bath Hamiltonian derivation procedure to molecular chromium bromide. Here, we have shown that the derived effective spin-bath models are good descriptions of the system’s low-energy spin excitations. Future work will address more complex materials, where the Schrieffer-Wolff transformation becomes essential to obtain good effective Hamiltonian descriptions for the low-energy spin dynamics of the material.
@@
The numerical data can be obtained from the authors on request. The spin finding procedure and the implemented Schrieffer Wolff transformation can be tested on https://cloud.quantumsimulations.de/ using the HQS Spin Mapper module.
@@ Appendix B Symmetry-specification of fermionic creation and annihilation operators
where we have fixed the order $\uparrow$,\,$\downarrow$ , ↓ for the fermionic sign. The fermionic annihilation operator $c_{\downarrow}:\mathcal{H}_{\textrm{local}}\rightarrow\mathcal{H}_{\mathrm{local}}$ for fermions with spin $s_{z}=-\frac{1}{2}$ acts on the vectors in the local Hilbert space and performs the linear map
@@
[…]

### v3

@@ III.5 Molecular Chromium bromide
We observe that the ground state energies of the effective Hamiltonians $\mathcal{P}H\mathcal{P}$, $\tilde{H}$, and $\mathcal{P}\tilde{H}\mathcal{P}$ are within a fraction of a Hartree from the true ground state energy $E_{0}^{H}$ of the original fermionic Hamiltonian $H$. It is noteworthy that the ground state energy $E_{0}^{\mathcal{P}\tilde{H}\mathcal{P}}$ of the effective spin-bath Hamiltonian $\mathcal{P}\tilde{H}\mathcal{P}$ deviates from the true ground state energy only on the order of $\mu\mathrm{Hartree}$. The respective energy differences amount to
$\Delta E_{0}^{\mathcal{P}\tilde{H}\mathcal{P}}=$ $\,\left(E_{0}^{H}-E_{0}^{\mathcal{P}\tilde{H}\mathcal{P}}\right)\simeq 2\times 10^{-5}\,\mathrm{Hartree}$
$<$ $\,\left(E_{0}^{H}-E_{0}^{\tilde{H}}\right)\simeq 2\times 10^{-3}\,\mathrm{Hartree}$
$<$ $\,\left(E_{0}^{H}-E_{0}^{\mathcal{P}H\mathcal{P}}\right)\simeq 8\times 10^{-3}\,\mathrm{Hartree}\,.$
In contrast, we find a significantly larger energy discrepancy $\Delta E_{0}^{\mathcal{P}_{0}H\mathcal{P}_{0}}=(E_{0}^{H}-E_{0}^{\mathcal{P}_{0}H\mathcal{P}_{0}})\simeq 2\times 10^{-1}\,\mathrm{Hartree}$ for the naive effective spin-bath Hamiltonian $\mathcal{P}_{0}H\mathcal{P}_{0}$.
@@
At first glance, we observe significant energy differences between the original fermionic Hamiltonian $H$ and the effective spin-bath Hamiltonians $\mathcal{P}H\mathcal{P}$ and $\mathcal{P}\tilde{H}\mathcal{P}$ for the excited states, while the transformed fermionic Hamiltonian $\tilde{H}$ appears to remain a reasonable approximation to the original Hamiltonian $H$ with $|\delta E_{n}|=|E_{n}^{H}-E_{n}^{\tilde{H}}|\leq 3\times 10^{-3}\,\mathrm{Hartree}$. We find that the $n=9$ lowest excited states correspond to an average local parity $\sum_{q\in\mathcal{S}}\langle P_{q}\rangle_{n}\geq-1$. This implies the presence of either no spin degree freedom or at most one distinct local spin degree of freedom instead of a total of three distinct ones, such that the effective three-spin-system Hamiltonians $\mathcal{P}H\mathcal{P}$ and $\mathcal{P}\tilde{H}\mathcal{P}$ are not sensible descriptions of the physics of the excited states with significant charge fluctuations. If we reduce the spectrum to eigenstates of $H$ with three spin-like orbitals, i.e. total local parity $\sum_{q\in{\mathcal{S}}}\langle P_{q}\rangle_{n_{\mathcal{S}}}\simeq-3$, a different picture emerges. This reduced spectrum is shown in Fig. 9 (a). The corresponding energy discrepancies $|\delta E_{n_{\mathcal{S}}}|$ from the true eigen energies of $H$ are shown in Fig. 9 (b). The effective Hamiltonians $\mathcal{P}H\mathcal{P}$ and $\mathcal{P}\tilde{H}\mathcal{P}$ appear to capture this reduced spectrum well. We observe a maximum energy discrepancy $|\delta E_{n_{\mathcal{S}}}|=1.5\times 10^{-3}\,\mathrm{Hartree}$ for the effective transformed Hamiltonian $\mathcal{P}\tilde{H}\mathcal{P}$, while we see a larger discrepancy $|\delta E_{n_{\mathcal{S}}}|=8\times 10^{-3}\,\mathrm{Hartree}$ for the effective Hamiltonian $\mathcal{P}H\mathcal{P}$. The effective Hamiltonian $\mathcal{P}_{0}H\mathcal{P}_{0}$ remains a bad approximation throughout this reduced energy spectrum. This highlights the necessity of a suitable orbital basis for the derivation of effective Hamiltonians describing the spin physics. In the parity-optimized orbital basis, for the given example of $\mathrm{Cr}\mathrm{Br}^{3-}_{6}$, the terms coupling to the charge degree of freedom of the electron densities in the spin-like orbitals $\phi_{q\in\mathcal{S}}$ are suppressed to the extent that the simple effective Hamiltonian $\mathcal{P}H\mathcal{P}$ already becomes a relatively good description of the spin excitation spectrum. The subsequent Schrieffer-Wolff transformation and utilization of the corresponding effective transformed Hamiltonian $\mathcal{P}\tilde{H}\mathcal{P}$ then yield smaller additional improvements. This emphasizes the importance of using a suitable single particle basis. In constrast, we observe no such drastic suppression of the undesirable coupling terms for the basis of the CASSCF canonical molecular orbitals.
@@
We have shown that our optimization procedure is able to correctly identify the $N_{\phi_{q}}=3$ spin-like orbitals such that a subset of the energy spectrum of $\mathrm{Cr}\mathrm{Br}^{3-}_{6}$ can be accurately reproduced by a system of three spin degrees of freedom coupled to a fermionic environment. We further observe that our Schrieffer-Wolff transformation approach yields a transformed Hamiltonian $\tilde{H}$ that, if projected to the Hilbert space representing three spins in a fermionic environment, improves upon the description given by the original Hamiltonian $H$ projected to the same subspace. As shown in Fig. 9 the energy of the ground state and the low energy spin excited states are much closer to the original model if we apply our Schrieffer-Wolff method before projecting the spin-like orbitals on pure spin orbitals. We validate that the local parity of the spin-like basis orbitals can be regarded as a good indicator of the accuracy of the effective spin-bath Hamiltonian description $\mathcal{P}\tilde{H}\mathcal{P}$ derived with our method.
@@ IV Conclusion
Effective model Hamiltonians provide enormous value for the investigation of individual aspects of complex systems. Here, we have presented a procedure for the automatic derivation of model Hamiltonians of coupled spin degrees of freedom embedded in a fermionic environment. These Hamiltonians provide a means to enhance the study of the dominant low temperature spin physics of complex real materials. We have shown that to investigate the spin physics of such materials properly, one first needs to operate within a suitable single particle orbital basis, where a subset of the basis orbitals become realizations of spin degrees of freedom at low energies. We have introduced the local parity as a sensible metric for the spin-like behavior of electron densities in the basis orbitals. We have proven that the natural orbital basis does not generally correspond to the basis that features the spin-like orbitals. We have provided an example toy model for which the natural orbital basis is significantly worse than the initial basis for the study of spin physics. We also studied a well-known singlet-diradical molecule, namely closed configuration para-benzyne, and demonstrated that the two well-known non-trivial spin-like orbitals are correctly determined by our local parity optimization procedure. We find that the optimization of the local parity yields an orbital basis that typically results in a many particle basis of the Hilbert space with small coupling and significant energy gap between the different subspaces. Our extended Schrieffer-Wolff transformation method allows to also transform generic Hamiltonians with complicated interaction terms beyond the density-density interactions of the toy models that the Schrieffer-Wolff transformation has so far mostly been applied to. There is also no need for the block-diagonal part $H_{0}$ to be easily diagonalizable, because no calculation of the spectrum of $H_{0}$ is required in Schrieffer-Wolff transformation our method. We have shown that our Schrieffer-Wolff transformation method reproduces established results for the well-known single impurity Anderson model and the disordered Fermi-Hubbard model Hamiltonians. For these models we corroborate that the local parity of the spin-like orbitals is a good predictor for the quality of the effective spin-bath Hamiltonian as the description of the low energy spin dynamics. Lastly, we have applied the full spin-bath Hamiltonian derivation procedure to molecular chromium bromide. Here, we have shown that the derived effective spin-bath models are good descriptions of the system’s low-energy spin excitations. We would like to point out that while we used an approximate multi reference method to determine the spin-like orbitals, the Schrieffer–Wolff mapping bypasses the standard quantum chemistry methods directly projecting the quantum chemistry model on the spin / spin-bath model. Being able to perform these calculations for quantum chemistry descriptions including all four index terms is provided by our representation of the Schrieffer-Wolff equation as a linear set of equations. Future work will address more complex materials, where the Schrieffer-Wolff transformation becomes essential to obtain good effective Hamiltonian descriptions for the low-energy spin dynamics of the material.
@@
The numerical data can be obtained from the authors on request. The spin finding procedure and the imple- mented Schrieffer Wolff transformation can be tested on https://cloud.quantumsimulations.de/ using the HQS Spin Mapper module.
@@ Appendix B Symmetry-specification of fermionic creation and annihilation operators
where we have fixed the order $\uparrow,\,\downarrow$ for the fermionic sign. The fermionic annihilation operator $c_{\downarrow}:\mathcal{H}_{\textrm{local}}\rightarrow\mathcal{H}_{\mathrm{local}}$ for fermions with spin $s_{z}=-\frac{1}{2}$ acts on the vectors in the local Hilbert space and performs the linear map
@@
where $\mathbf{0}_{\mathcal{H}}$ denotes the zero element of the Hilbert space. We see that $c_{\downarrow}$ maps two of the states, namely $|0\rangle$ and $|\uparrow\rangle$, map to the zero element. There are thus two non-vanishing components of $c_{\downarrow}$ which map $|\downarrow\rangle\rightarrow|0\rangle$ and $|\uparrow\downarrow\rangle\rightarrow|\uparrow\rangle$. We can choose to distinguish the blocks of the Hilbert space based on their total particle number $n=\sum_{\sigma}c^{\dagger}_{\sigma}c_{\sigma}$. We consequently identify the state $|\uparrow\downarrow\rangle$ as belonging to the block characterized by the eigenvalue $\Lambda_{q}=n=2$ and state $|\downarrow\rangle$ to the block characterized by the eigenvalue $\Lambda_{q}=n=1$. Given the choice of $n$ as the operator distinguishing the blocks of the Hilbert space, an intuitive choice for the operator $A$ is the local number of fermions with opposite spin $s_{z}=\frac{1}{2}$, i.e. the operator $n_{\uparrow}=c^{\dagger}_{\uparrow}c_{\uparrow}$. It is easily seen that the choice $A=n_{\uparrow}$ satisfies